An arc with radius R=0.25m and length  l=π/12 meter carries 3A current. What is the magnetic field B at the center of the arc.


1
Expert's answer
2021-01-14T16:35:11-0500

Magnetic field of any arc:

"B = \frac{μ_0}{4π} \times \frac{I}{R} \times α \\

α = \frac{AB}{R} \\

B = \frac{μ_0}{4π} \times \frac{3}{0.25} \times \frac{π}{3} = μ_0 \;Tesla"

Answer: μ0